# SoK: Exploring the Potential of Large Language Models for Improving Digital Forensic Investigation Efficiency

### v1

@@ Abstract.
The growing number of cases requiring digital forensic analysis raises concerns about law enforcement’s ability to conduct investigations promptly. Consequently, this systemisation of knowledge paper delves into the potential and effectiveness of integrating Large Language Models (LLMs) into digital forensic investigation to address these challenges. A thorough literature review is undertaken, encompassing existing digital forensic models, tools, LLMs, deep learning techniques, and the utilisation of LLMs in investigations. The review identifies current challenges within existing digital forensic processes and explores both the obstacles and possibilities of incorporating LLMs. In conclusion, the study asserts that the adoption of LLMs in digital forensics, with appropriate constraints, holds the potential to enhance investigation efficiency, improve traceability, and alleviate technical and judicial barriers faced by law enforcement entities.
@@
Digital Forensics, Large Language Models, Investigative Process, Challenges
journalyear: 2024copyright: acmlicensedconference: The 19th International ARES Conference on Availability, Reliability and Security; July 30–August 2, 2024; Vienna, Austriabooktitle: The 19th International ARES Conference on Availability, Reliability and Security, July 30–August 2, 2024, Vienna, Austriadoi: https://doi.org/10.1145/3407023.3407068isbn: 978-1-4503-8833-7/20/08copyright: noneconference:  ;  ;
1. Introduction
@@
With the pervasive growth of Information and Communication Technology (ICT) and information systems, cybercrimes have witnessed a significant surge in recent years (Ali, 2019). As a further compounding factor, the number of “traditional” police investigations including digital evidence is also ever-increasing (Du and Scanlon, 2019). Addressing and investigating this volume of cases presents substantial challenges.
@@
Artificial General Intelligence (AGI) and Large Language Models (LLMs) have become prominent topics of global discussion, prompting researchers to intensify their investigations by leveraging the capabilities of LLMs. The usage of LLMs within the scientific community experienced a rapid surge after 2022, notably with the advent of OpenAI’s ChatGPT platform. In a remarkably short period, this topic has garnered attention from academia, industry, and the research community at large (Ray, 2023). Simultaneously, researchers are exploring the potential of LLMs across various domains and assessing their impact on the future of science and society. This inquiry also includes an examination of the potential harmfulness associated with the deployment of LLMs (Brown et al., 2020; Nozza et al., 2022). In other words, the use of LLMs in various tasks can be a double-edged sword, necessitating careful consideration depending on the specific situations and contexts.
@@
Given the swiftly evolving landscape of LLMs, it is prudent to look into various types and their unique capabilities. A nuanced understanding of the strengths and characteristics of different LLMs can contribute to more informed and effective applications within the dynamic field of digital forensics (DF).
@@
The main objective of this article is to explore the viability of integrating LLMs into the fundamental phases of the modern DF process. This exploration involved a meticulous and systematic review, adhering to the standards set by Kitchenham (2004). The literature selected for this review was exclusively in English, and non-article sources such as books were excluded. The focus of the review encompassed empirical studies, case studies, and literature reviews within the domain of DF investigation methods, models, frameworks, efficiency, and challenges, as well as LLM capabilities, challenges, and architectures. Additionally, the study underscores the utilisation of autonomous agents based on LLMs, emphasising their capabilities and the potential efficiency gains achievable in the field of DF. In light of the fast-paced advancements and recent explosion in LLM-focused research, a substantial influx of LLM-focused research papers has occurred since the launch of ChatGPT in late 2022. Due to this fast pace, a large number of research articles exist solely as preprints on preprint services, e.g., arXiv or OpenReview. To give two examples, the initial papers for GPT-4 (Achiam et al., 2023) and LLaMA (Touvron et al., 2023) are both only published on arXiv, but have garnered thousands of citations each. Despite their preprint status, these papers offer essential insights critical for contemporary research and dialogue within the domain, making their incorporation into this paper necessary to provide the most up-to-date knowledge and perspectives.
@@
The paper is structured as follows: Section 2 provides a comprehensive background for the review, delving into existing DF process models, the challenges inherent in DF, and a detailed overview of the current work conducted with the utilisation of LLMs within DF. In Section 3 the paper delves into the realm of Natural Language Processing (NLP), elucidating the working principles of LLMs, their architectural foundations, and the specifics surrounding specially trained LLMs. Section 4 provides an in-depth review focusing on the capabilities and benchmark information of LLMs trained for coding tasks, as well as those tailored for vision assistance. Section 4 explores the synergy between DF and LLMs, detailing how LLMs can be effectively employed in each phase of the DF process model. Finally, in Section 6, the paper summarises the future challenges associated with integrating LLMs with automated agents within the DF domain. It outlines potential avenues for future research and development, shedding light on the path of future DF investigations employing LLMs. The discussion encompasses not only the potential negative impacts but also the practical difficulties and risks in real-world environments.
2. Background
@@
DF is a process to identify, preserve, analyse, and document digitally recorded data, which originate in electronic devices such as computers, servers, smartphones, and IoT devices (Baryamureeba and Tushabe, 2004). This exercise is required in most criminal cases. Data collected in this process are kept unchanged and safe to present in a court case or to support future investigations conducted by law enforcement agencies (Mukherjee and Haque, 2018).
2.1. Digital Forensic Process Models
@@
DF process models consist of a series of activities that help to standardise the investigative process (Du et al., 2017) and outline the phases; collection, preservation of evidence, examination or analysis, and reporting.
Figure 1. Traditional digital forensic process model (Casey, 2011)
@@
DF encompasses various subdisciplines such as computer forensics, mobile device forensics, memory forensics, network forensics, and cloud forensics, each employing distinct processes reflected in a plethora of models within the literature (Scanlon et al., 2023a; Wu et al., 2020). These models often share phases but differ in their focus and execution. For instance, Al-Dhaqm et al. (2020) proposed a mobile forensic model that adds a preparatory phase and bifurcates the analysis stage into examination and analysis phases. To accommodate the complexities of computer, network, cloud, and smart device forensics, Lutui (2016) introduced a multidisciplinary model that necessitates diverse skills for effective investigation, covering incident detection to evidence storage.
Casey’s model, as shown in Figure 1, includes phases like incident recognition, evidence collection, preservation, and presentation, with the examination phase detailed into recovery, harvesting, reduction, and classification (Casey, 2011). This model emphasises the critical nature of maintaining evidence integrity and the requirement for expert analysis to extract and interpret pertinent information, culminating in a report suitable for legal scrutiny (Casey, 2011). Notably, the analysis or examination stage is pivotal in all models, demanding specialised knowledge in the relevant DF area (Mir et al., 2016; Prayudi et al., 2020).
@@
The advent of cloud computing has led to the Digital Forensic as a Service (DFaaS) model by van Baar et al., integrating evidence preservation and analysis into an automated, secure software service, marking a significant evolution in forensic methodologies (van Beek et al., 2015; Du et al., 2017; van Beek et al., 2020).
2.2. Existing Challenges in Digital Forensics
@@
DF is an evolving field, yet the literature highlights that it still undergoes changes to address ongoing challenges and advancements. Dubey et al. (2023) assert that DF faces key challenges, including the complexity of data and its volume, a lack of standardisation, inadequacies in the power of existing tools to support investigations and issues related to timelines.
@@
In addition to the previously mentioned challenges, issues such as scope creep in cases due to complexity and vast data, as well as the critical tasks of selecting and prioritising the right set of evidence, and efficiently allocating time and investigators for the chosen evidence, were highlighted (Kalaimannan et al., 2013). Koper et al. (2014) focus on certain issues from the investigator’s perspective, including challenges in adapting to a system, unexpected time-saving negations, and frustrations among officers arising from timelines and the adoption of complex systems.
@@
Automating the DF process using existing technology appears to be a promising solution for addressing issues related to time management and effectiveness (Michelet et al., 2023). However, an ongoing challenge revolves around measuring the accuracy of investigations and ensuring the verification of the automated process. This aspect remains an open area that requires further attention and resolution (James and Gladyshev, 2013).
2.3. Existing Work With LLMs in Digital Forensics
@@
Scanlon et al. (2023a) analysed using ChatGPT for DF. In their assessment, the authors evaluated the programming, incident narration, keyword list creation, and DF teaching abilities of ChatGPT. Their conclusion highlighted that while ChatGPT exhibited some hallucinations in the output results, it still serves as an effective assistant for code generation.
@@
Timeline reconstruction aids investigators in deducing the behaviour of an event. In line with timeline regeneration, Silalahi et al. (2023) proposed a method to detect anomalies in a drone flight by employing sentiment analysis with the assistance of a pre-trained LLM. Their approach successfully discerned differences between normal and abnormal events with an accuracy of 92.5%.
@@
Hansken, a DFaaS platform created by the Netherlands Forensic Institute, is designed to assist investigators in handling evidence and conducting investigations more efficiently (van Beek et al., 2015). ChatGPT has been utilised as an assistance for the Hansken DFaaS system, contributing to streamlined processes and improved support for investigators (Henseler and van Beek, 2023). In a successful experiment, evidence traces were provided to ChatGPT, and it was tasked with analysing the case, ultimately providing conclusions. This utilisation of ChatGPT showcases its potential in assisting with the analytical aspects of investigations, highlighting its ability to process and interpret evidence data (Henseler and van Beek, 2023).
3. Large Language Models
@@
This section explores LLMs, concentrating on three principal aspects. Initially, it explores the architecture of LLMs, detailing their design and function. Next, it assesses the usability of LLMs, underscoring the features and capabilities that render them apt for a wide range of tasks. Finally, it showcases the versatility of LLMs by discussing their application across various fields, demonstrating their wide-reaching impact and the extensive scope of their applications.
3.1. Natural Language Processing
@@
Popular LLMs such as Generative Pre-trained Transformer (GPT) (Dale, 2021), Language Model for Dialogue Applications (LaMDA) (Thoppilan et al., 2022), Pathways Language Model (PaLM) (Chowdhery et al., 2023), Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019), and Large Language Model Meta AI (LLaMA) (Touvron et al., 2023) stem from advancements in Natural Language Processing (NLP). NLP, focusing on language-based tasks, utilises both traditional and deep learning models to enable applications such as language translation, text processing, and speech recognition (Reshamwala et al., 2013).
@@
Deep learning, a branch of machine learning, uses complex computational layers and adaptive weights to enhance prediction accuracy, offering a more refined analysis than conventional machine learning (LeCun et al., 2015). It has excelled in image and speech recognition, and natural language understanding, mimicking the decision-making process of the human brain through artificial neurons. These neurons form networks capable of intricate pattern recognition and data analysis. Central to deep learning are neural networks with multiple hidden layers, which autonomously learn and feature-extract from data, bypassing the need for manual variable selection. This automatic feature extraction makes them exceptionally adept at handling complex tasks (Wang et al., 2017).
3.2. LLMs
@@
An LLM is a language model employing neural networks with billions of parameters, trained on extensive text data. These models are engineered to comprehend and generate human language. Fundamentally, they rely on multiple neural network architectures, enabling them to recognise the relationships between words and phrases within sentences (Fan et al., 2023; Shen et al., 2023a). These architectures have been a transformative force in the field of natural language processing. Its capability to excel across a diverse array of language-related tasks distinguishes it as a game-changer, in contrast to being tailored for a singular, specific task.
3.3. Architecture of LLMs
@@
LLMs utilise deep learning, particularly neural networks, to process and produce human language. Fundamentally, a language model operates with letters or words, but since machine learning algorithms and neural networks require vector inputs, words are vectorised. Each word in the vocabulary is assigned a unique numerical value for input into neural networks. Through initial random weight assignments and subsequent backpropagation, words acquire numerical positions reflecting their semantic similarity, culminating in a word embedding model (Lai et al., 2016).
3.3.1. Word to Vectors

### v2

# Exploring the Potential of Large Language Models for Improving Digital Forensic Investigation Efficiency

@@ Abstract
The growing number of cases that require digital forensic analysis raises concerns about the ability of law enforcement to conduct investigations promptly. Consequently, this paper delves into the potential and effectiveness of integrating Large Language Models (LLMs) into digital forensic investigation to address these challenges. A comprehensive literature review is carried out, encompassing existing digital forensic models, tools, LLMs, deep learning techniques, and the use of LLMs in investigations. The review identifies current challenges within existing digital forensic processes and explores both the obstacles and possibilities of incorporating LLMs. In conclusion, the study asserts that the adoption of LLMs in digital forensics, with appropriate constraints, has the potential to improve investigation efficiency, improve traceability, and alleviate technical and judicial barriers faced by law enforcement entities.
keywords:
@@
Digital Forensics , Large Language Models , LLM , Investigative Process , Challenges
journal: Elsevier
[…]
[label1]organization=Forensics and Security Research Group, School of Computer Science, addressline=University College Dublin, city=Belfield, state=Dublin 4, country=Ireland
[…]
[label2]organization=Faculty of Law, Criminal Justice and Public Administration, addressline=University of Lausanne, city=Lausanne, country=Switzerland
1 Introduction
@@
With the pervasive growth of Information and Communication Technology (ICT) and information systems, cybercrimes have seen a significant increase in recent years [1]. As a further compounding factor, the number of “traditional” police investigations that include digital evidence is also increasing [2]. Addressing and investigating this volume of cases presents substantial challenges.
@@
Artificial General Intelligence (AGI) and Large Language Models (LLMs) have become prominent topics of global discussion, prompting researchers to intensify their investigations by leveraging the capabilities of LLMs. The usage of LLMs within the scientific community experienced a rapid surge after 2022, notably with the advent of OpenAI’s ChatGPT platform. In a remarkably short period, this topic has garnered attention from academia, industry, and the research community at large [3]. Simultaneously, researchers are exploring the potential of LLMs across various domains and assessing their impact on the future of science and society. This inquiry also includes an examination of the potential harmfulness associated with the deployment of LLMs [4, 5]. In other words, the use of LLMs in various tasks can be a double-edged sword, necessitating careful consideration depending on the specific situations and contexts.
@@
Given the rapidly evolving landscape of LLMs, it is prudent to look into various types and their unique capabilities. A nuanced understanding of the strengths and characteristics of different LLMs can contribute to more informed and effective applications within the dynamic field of digital forensics (DF).
@@
The main objective of this article is to explore the viability of integrating LLMs into the fundamental phases of the modern DF process. This exploration involved a meticulous and systematic review, adhering to the standards set by Kitchenham [6]. The literature selected for this review was exclusively in English, and non-article sources such as books were excluded. The focus of the review included empirical studies, case studies, and literature reviews within the domain of DF investigation methods, models, frameworks, efficiency, and challenges, as well as LLM capabilities, challenges, and architectures. Additionally, the study underscores the utilisation of autonomous agents based on LLMs, emphasising their capabilities and the potential efficiency gains achievable in the field of DF.
As per the procedure outlined by Kitchenham [6], the quality of the chosen literature was assessed. Design, analysis, and conclusion served as the primary criteria for determining each paper’s quality. In the design benchmark, attention was given to the clarity of the paper’s objectives, definition of variables and measurements, and transparency of methodology. Regarding the analysis benchmark, focus was placed on the description and justification of analytical methods, clarity of the scoring system, and coherence of the analysis. The conclusion benchmark was particularly significant – involving a thorough examination to ensure that the paper’s conclusion adequately addressed study questions, discussed key findings and limitations, and acknowledged any negative aspects.
In light of fast-paced advancements and the recent explosion in LLM-focused research, a substantial influx of LLM-focused research papers has occurred since the launch of ChatGPT in late 2022. Due to this fast pace, many research articles exist solely as preprints on preprint services, e.g., arXiv or OpenReview. To give two examples, the initial papers for GPT-4 [7] and LLaMA [8] are only published on arXiv, but have garnered thousands of citations each. Despite their preprint status, these articles offer essential insights critical for contemporary research and dialogue within the domain, making their incorporation into this article necessary to provide the most up-to-date knowledge and perspectives.
@@
The paper is structured as follows: Section 1.1 provides a comprehensive background for the review, delving into existing DF process models, the challenges inherent in DF, and a detailed overview of the current work conducted with the utilisation of LLMs within DF. In Section 2 the paper delves into the realm of Natural Language Processing (NLP), elucidating the working principles of LLMs, their architectural foundations, and the specifics surrounding specially trained LLMs. Section 3 provides an in-depth review that focuses on the capabilities and benchmark information of LLMs trained for coding tasks, as well as those tailored for vision assistance. Section 3 explores the synergy between DF and LLMs, detailing how LLMs can be effectively employed in each phase of the DF process model. Finally, in Section 6, the paper summarises the future challenges associated with integrating LLMs with automated agents within the DF domain. The conclusion outlines potential avenues for future research and development, shedding light on the path of future DF investigations employing LLMs. The discussion covers not only the potential negative impacts but also the practical difficulties and risks in real world environments.
1.1 Digital Forensic Context
@@
DF is a process for identifying, preserving, analysing, and documenting digitally recorded data, which originate in electronic devices such as computers, servers, smartphones, and IoT devices [9]. This exercise is required in most criminal cases. Data collected in this process are kept unchanged and safe to present in a court case or to support future investigations conducted by law enforcement agencies [10].
1.1.1 Digital Forensic Process Models
@@
DF process models consist of a series of activities that help to standardise the investigative process [11] and outline the phases; collection, preservation of evidence, examination or analysis, and reporting.
Figure 1: Traditional digital forensic process model [12]
@@
DF encompasses various subdisciplines such as computer forensics, mobile device forensics, memory forensics, network forensics, and cloud forensics, each employing distinct processes reflected in a plethora of models within the literature [13, 14]. These models often share phases but differ in their focus and execution. For example, Al-Dhaqm et al. [15] proposed a mobile forensic model that adds a preparatory phase and bifurcates the analysis stage into examination and analysis phases. To accommodate the complexities of computer, network, cloud, and smart device forensics, Lutui [16] introduced a multidisciplinary model that requires diverse skills for effective investigation, ranging from incident detection to evidence storage.
Casey’s model, as shown in Figure 1, includes phases such as incident recognition, evidence collection, preservation, and presentation, with the examination phase detailed in recovery, harvesting, reduction, and classification [12]. During incident recognition, the focus is on identifying the incident itself, potential evidentiary sources, and expected digital evidence types, as well as delineating the scope of the ensuing investigation. Conversely, investigators systematically acquire pertinent evidence from various sources encompassing computers, smartphones, storage media, and networks during collection and seizure. Preservation assumes paramount importance in upholding the integrity of the evidence, necessitating specific and accurate measures to ensure the unmodified condition of collected data throughout the investigative process. The overarching objective remains the meticulous safeguarding of evidence integrity.
The subsequent phase entails examination, wherein analysts rigorously scrutinise the gathered data to extract pertinent information. This endeavour may involve the employment of diverse forensic hardware and software tools and techniques. The examination process encompasses interpreting extracted information to draw conclusive inferences regarding the events under scrutiny. This phase frequently demands a profound comprehension of both the technology employed and the context surrounding the evidence.
Next is the reporting phase, where the findings derived from the analysis are systematically presented in a format suitable for legal adjudication. This may involve preparing comprehensive reports and providing expert testimony in a court of law. This model emphasises the critical nature of maintaining evidence integrity and provenance, and the requirement for expert analysis to extract and interpret pertinent information. This culminates in a report suitable for legal scrutiny. In particular, the analysis or examination stage is crucial in all models, demanding specialised knowledge in the relevant DF area [17, 18].
@@
The advent of cloud computing has led to the Digital Forensic as a Service (DFaaS) model by van Baar et al., which integrates evidence preservation and analysis into an automated and secure software service, marking a significant evolution in forensic methodologies [20, 11, 21].
1.1.2 Existing Challenges in Digital Forensics
@@
DF is an evolving field, yet the literature highlights that it still undergoes changes to address ongoing challenges and advancements. Dubey et al. [22] assert that DF faces key challenges, including the complexity of data and its volume, a lack of standardisation, inadequacies in the power of existing tools to support investigations and issues related to timelines.
@@
In addition to the previously mentioned challenges, other issues persist including scope creep in cases due to complexity and the vast data involved, selecting and prioritising the right set of evidence, and efficiently allocating time and investigators for the chosen evidence [23]. Koper et al. [24] focus on a number of these issues from the investigator’s perspective, including challenges in adapting to a system, unexpected time sinks, and frustrations among officers arising from expected operational timelines and the adoption of complex systems. The contemporary landscape of forensic science is characterised by a notable deficiency in proficient agents, exacerbated by the swiftly evolving standards, practices, tools, and techniques within the field. Moreover, the predominant emphasis of law enforcement roles on fieldwork, as opposed to dedicated DF duties, has further contributed to the scarcity of adept human expertise in this domain [25].
@@
Automating the DF process using existing technology appears to be a promising solution to address issues related to time management and effectiveness [26]. However, an ongoing challenge revolves around measuring the accuracy of investigations and ensuring the verification of the automated process. This aspect remains an open area that requires more attention and resolution [27].
1.1.3 Existing Work With LLMs in Digital Forensics
@@
Scanlon et al. [13] analysed using ChatGPT for DF. In their assessment, the authors evaluated the programming, incident narration, keyword list creation, and DF teaching abilities of ChatGPT. Their conclusion highlighted that while ChatGPT exhibited some hallucinations in the output results, it still serves as an effective assistant for code generation.
@@
Timeline reconstruction helps investigators deduce the chronological “story” of an event. In line with timeline regeneration, Silalahi et al. [28] proposed a method to detect anomalies in a drone flight by employing sentiment analysis with the assistance of a pre-trained LLM. Their approach successfully discerned differences between normal and abnormal events with an accuracy of 92.5%.
@@
Hansken, a DFaaS platform created by the Netherlands Forensic Institute, is designed to help investigators handle evidence and conduct investigations more efficiently [29]. ChatGPT has been used as an assistance for the Hansken DFaaS system, contributing to streamlined processes and improved support for investigators [30]. In a successful experiment, evidence traces were provided to ChatGPT, and it was tasked with analysing the case, ultimately providing conclusions. This utilisation of ChatGPT showcases its potential in assisting with the analytical aspects of investigations, highlighting its ability to process and interpret evidence data [30].
2 Background of Large Language Models
@@
This section explores LLMs, concentrating on three principal aspects. Initially, it explores the architecture of LLMs, detailing their design and function. Next, it assesses the usability of LLMs, underscoring the features and capabilities that render them apt for a wide range of tasks. Finally, it showcases the versatility of LLMs by discussing their applications across various fields, demonstrating their wide-reaching impact and the extensive scope of their applications.
2.1 Natural Language Processing
@@
Popular LLMs such as Generative Pre-trained Transformer (GPT) [31], Language Model for Dialogue Applications (LaMDA) [32], Pathways Language Model (PaLM) [33], Bidirectional Encoder Representations from Transformers (BERT) [34], and Large Language Model Meta AI (LLaMA) [8] stem from advances in Natural Language Processing (NLP). NLP, focusing on language-based tasks, utilises both traditional and deep learning models to enable applications such as language translation, text processing, and speech recognition [35].
@@
Deep learning, a branch of machine learning, uses complex computational layers and adaptive weights to improve prediction accuracy, offering a more refined analysis than conventional machine learning [36]. It has excelled in image and speech recognition, and natural language understanding, mimicking the decision-making process of the human brain through artificial neurons. These neurons form networks capable of intricate pattern recognition and data analysis. Central to deep learning are neural networks with multiple hidden layers, which autonomously learn and feature-extract from data, bypassing the need for manual variable selection. This automatic feature extraction makes them exceptionally adept at handling complex tasks [37].
2.2 LLMs
@@
An LLM is a language model employing neural networks with billions of parameters, trained on extensive text data. These models are engineered to comprehend and generate human language. Fundamentally, they rely on multiple neural network architectures, enabling them to recognise the relationships between words and phrases within sentences [38, 39]. These architectures have been a transformative force in natural language processing. Its capability to excel across a diverse array of language-related tasks distinguishes it as a game-changer, in contrast to being tailored for a singular, specific task.
2.3 Architecture of LLMs
@@
LLMs utilise deep learning, particularly neural networks, to process and produce human language. Fundamentally, a language model operates with letters or words, but since machine learning algorithms and neural networks require vector inputs, words are vectorised. Each word in the vocabulary is assigned a unique numerical value for input into neural networks. Through initial random weight assignments and subsequent backpropagation, words acquire numerical positions reflecting their semantic similarity, culminating in a word embedding model [40].
2.3.1 Word to Vectors